### v2

# A note on one-variable theorems for NSOP

Abstract
We give an example of an SOP theory $T$, such that any $L(M)$-formula $\varphi(x,y)$ with $|y|=1$ is NSOP. We show that any such $T$ must have the independence property. We also give a simplified proof of Lachlan’s [3] theorem that if every $L$-formula $\varphi(x,y)$ with $|x|=1$ is NSOP, then $T$ is NSOP.
[…]
2020 Mathematical Subject Classification: 03C45.
Key words and phrases: NSOP
1 Introduction
Fix a complete theory $T$. Recall that a formula $\varphi(x,y)$ has the strict order property (SOP) if in some model $M$ there is a sequence $b_{0},b_{1},\ldots$ with
$\varphi(M,b_{0})\subsetneq\varphi(M,b_{1})\subsetneq\cdots$
An $L$-theory has the SOP iff some $L$-formula has it. A formula or theory is NSOP if it doesn’t have the SOP. A classic theorem of Lachlan [3] shows that NSOP can be checked on one-variable formulas:
Theorem 1.1 (Lachlan).
If $T$ has the SOP, then there is an $L$-formula $\varphi(x,y)$ with the SOP, with $|x|=1$.
Lachlan’s proof is short but rather convoluted, so in Section 2 we give what is hopefully a simpler proof.
Remark 1.2.
Lachlan’s result is analogous to Shelah’s one-variable theorems for stability and NIP [6, Theorems 2.13, 4.6], Chernikov’s one-variable theorem for NTP2 [1, Theorem 2.9, Lemma 3.2], and Nick Ramsey’s one-variable theorem for NSOP1 [5]. These theorems are identical to Theorem 1.1 with “SOP” replaced with “unstable”, “IP”, “TP2”, and “SOP1”, respectively.
Lachlan’s result suggests the following question:
Question 1.3.
If $T$ has the SOP, is there necessarily an SOP $L$-formula $\varphi(x,y)$ with $|y|=1$?
The analogous questions for NIP and stability have positive answers because stability and NIP are symmetric notions: a formula $\varphi$ is stable or NIP iff the opposite formula $\varphi^{opp}(y;x):=\varphi(x;y)$ is stable or NIP.
In contrast, NSOP is not symmetric, and the answer to Question 1.3 is no:
Example 1.4.
Let $T$ be the theory of the structure $(\mathbb{Q},B)$, where $B(x,y,z)$ means that either $x<y<z$ or $x>y>z$. Note that $\operatorname{tp}(a,b)\neq\operatorname{tp}(b,a)$ for any distinct singletons $a,b$, in any model of $T$. The formula $B(x;y,z)$ has the SOP. However, no $L$-formula $\varphi(x;y)$ with $|y|=1$ has the SOP. Otherwise, take an ascending chain
$\varphi(\mathbb{M};b_{0})\subsetneq\varphi(\mathbb{M};b_{1})\subsetneq\cdots$
Then $\operatorname{tp}(b_{0},b_{1})=\operatorname{tp}(b_{1},b_{0})$, so $\varphi(\mathbb{M};b_{1})\subsetneq\varphi(\mathbb{M};b_{0})$, which is absurd.
The circular order is another example, again because $\operatorname{tp}(a,b)=\operatorname{tp}(b,a)$ for any two distinct singletons $a,b$.
But what if we allow the formula to mention parameters from the model?
Question 1.5.
If $T$ has the SOP, and $\mathbb{M}$ is a monster model of $T$, is there necessarily an SOP $L(\mathbb{M})$-formula $\varphi(x,y)$ with $|y|=1$?
For instance, in Example 1.4, the formula $\varphi(y;z)\equiv B(0,y;z)$ has the strict order property. One striking result in this direction is due to Pierre Simon [8]:
Theorem 1.6 (Simon).
Let $T$ be a theory with monster model $\mathbb{M}$.
  1. 1.
    If $T$ is unstable, then some $L(\mathbb{M})$-formula $\varphi(x,y)$ with $|x|=|y|=1$ is unstable.
  2. 2.
    If $T$ has the independence property, then some $L(\mathbb{M})$-formula $\varphi(x,y)$ with $|x|=|y|=1$ has the independence property.
Could we expect the same to work for NSOP? Unfortunately, the answer is again no:
Theorem 1.7.
There is a theory $T$ with the SOP, such that every $L(M)$-formula $\varphi(x,y)$ with $|y|=1$ is NSOP.
Although the theory $T$ has a simple description (see Subsection 4.2), the proof of Theorem 1.7 is extremely unpleasant, taking up the bulk of this paper. I would love to find a simpler counterexample in the spirit of Example 1.4.
On the other hand, the answer to Question 1.5 is yes if we assume $T$ is NIP:
Theorem 1.8.
Let $T$ be an NIP theory with monster model $\mathbb{M}$. If $T$ has the SOP, then there is an $L(\mathbb{M})$-formula $\varphi(x,y)$ with the SOP with $|x|=|y|=1$.
The proof is quite easy, falling directly out of the standard proof that “stable $=$ NIP $+$ NSOP”. We give the details in Section 3. Theorem 1.8 shows that any theory $T$ as in Theorem 1.7 must have both the SOP and IP, which perhaps explains why $T$ must be complicated.
Theorem 1.7 can be recast more geometrically as a statement about definable posets. Recall that a poset $(P,\leq)$ has finite height if there is a finite upper bound $n$ on the cardinality of chains in $P$. Any formula $\varphi(x,y)$ determines a partial order on $\mathbb{M}^{y}$ in which
$b<b^{\prime}\iff\varphi(\mathbb{M},b)\subsetneq\varphi(\mathbb{M},b^{\prime}).$
The formula $\varphi(x,y)$ has the SOP if and only if this poset has infinite height. Theorem 1.7 is thus equivalent to the following:
Theorem 1.9.
There is an SOP theory $T$, such that any definable poset $(P,\leq)$ with $P\subseteq\mathbb{M}^{1}$ has finite height.
This can be contrasted with the situation with linear orders:
Fact 1.10 ([2, Lemma 5.4]).
If there is an infinite definable linear order $(P,\leq)$, then there is one with $P\subseteq\mathbb{M}^{1}$.
1.1 Conventions
Throughout, letters $a,b,c,\ldots,x,y,z,\ldots$ represent tuples of elements or variables, rather than single elements or variables. The length of a tuple $x$ is written $|x|$. If $x$ is a tuple of variables and $M$ is a structure, then $M^{x}$ is the set of tuples in $M$ of length $|x|$. By default we work in a monster model $\mathbb{M}$ of a complete $L$-theory $T$. If $A$ is a set of parameters, then $L(A)$ is the expansion of $L$ by naming the elements of $A$ as constants. We try to clarify whether “formula” means $L(\mathbb{M})$-formula or $L$-formula, whenever the difference matters. A partial order means a strict partial order. We write disjoint unions as $X\sqcup Y$.
2 A simplified proof of Lachlan’s theorem
An $L(\mathbb{M})$-formula $\varphi(x,y)$ is NSOP iff the poset of $\varphi$-sets $\{\varphi(\mathbb{M},b):b\in\mathbb{M}^{y}\}$ has finite height.
If $P$ is a poset with finite height, let $\operatorname{ht}:P\to\mathbb{N}$ be the height function. Thus $\operatorname{ht}(a)\geq k$ iff there is a chain $a_{0}<a_{1}<\cdots<a_{k}=a$. If an $L(\mathbb{M})$-formula $\varphi(x;y)$ is NSOP and $b\in\mathbb{M}^{y}$, let $\operatorname{ht}_{\varphi}(b)$ denote the height of $\varphi(\mathbb{M};b)$ in the poset of $\varphi$-sets. Then
$\varphi(\mathbb{M},b)\subsetneq\varphi(\mathbb{M},b^{\prime})\implies\operatorname{ht}_{\varphi}(b)<\operatorname{ht}_{\varphi}(b^{\prime}).$
Lemma 2.1.
If an $L(\mathbb{M})$-formula $\varphi(x,y;z)$ has the SOP, then some $L(\mathbb{M})$-formula $\psi(x;z)$ or $\theta(y;z)$ has the SOP.
Note that the $x,y,z$ appearing in $\psi(x;z)$ and $\theta(y;z)$ are the same $x,y,z$ appearing in $\varphi(x,y;z)$. In particular, the length of $z$ didn’t change.
Proof.
Suppose the lemma fails. Without loss of generality, $\varphi$ is an $L$-formula. For $a\in\mathbb{M}^{x}$, let $\varphi_{a}(y;z)$ be $\varphi(a,y,z)$. Then $\varphi_{a}$ is NSOP so $\operatorname{ht}_{\varphi_{a}}(c)$ makes sense for any $a\in\mathbb{M}^{x}$ and $c\in\mathbb{M}^{z}$. For $k\in\mathbb{N}$ let $\psi_{k}(x,z)$ be the $L$-formula such that
$\operatorname{ht}_{\varphi_{a}}(c)\geq k\iff\mathbb{M}\models\psi_{k}(a,c).$
Take an indiscernible sequence $c_{0},c_{1},c_{2},\ldots$ with
$\varphi(\mathbb{M},c_{0})\subsetneq\varphi(\mathbb{M},c_{1})\subsetneq\cdots$
For any $a\in\mathbb{M}^{x}$ and $k\in\mathbb{N}$ we have
$\varphi(a,\mathbb{M},c_{0})\subseteq\varphi(a,\mathbb{M},c_{1})$ ($\ast$)
$\operatorname{ht}_{\varphi_{a}}(c_{0})\leq\operatorname{ht}_{\varphi_{a}}(c_{1})$
$\operatorname{ht}_{\varphi_{a}}(c_{0})\geq k\implies\operatorname{ht}_{\varphi_{a}}(c_{1})\geq k$
$a\in\psi_{k}(\mathbb{M},c_{0})\implies a\in\psi_{k}(\mathbb{M},c_{1}),$
and so $\psi_{k}(\mathbb{M},c_{0})\subseteq\psi_{k}(\mathbb{M},c_{1})$. Equality must hold, or $\psi_{k}$ has the SOP by indiscernibility. Then for any $a\in\mathbb{M}^{x}$ and $k\in\mathbb{N}$, we have
$a\in\psi_{k}(\mathbb{M},c_{0})\iff a\in\psi_{k}(\mathbb{M},c_{1})$
$\operatorname{ht}_{\varphi_{a}}(c_{0})\geq k\iff\operatorname{ht}_{\varphi_{a}}(c_{1})\geq k$
$\operatorname{ht}_{\varphi_{a}}(c_{0})=\operatorname{ht}_{\varphi_{a}}(c_{1})$
$\varphi(a,\mathbb{M},c_{0})=\varphi(a,\mathbb{M},c_{1}),$
where the last line follows by ($\ast$). Thus $\varphi(\mathbb{M},c_{0})=\varphi(\mathbb{M},c_{1})$, a contradiction. ∎
Theorem 2.2 (Lachlan).
If an $L(\mathbb{M})$-formula $\varphi(x;y)$ has the SOP, then some $L(\mathbb{M})$-formula $\psi(z;y)$ has the SOP with $|z|=1$.
Note that the length of $y$ is the same in $\varphi$ and $\psi$.
Proof.
By induction on $|x|$ using Lemma 2.1. ∎
Theorem 2.3 (Lachlan).
If $T$ has the SOP, then some $L$-formula $\varphi(x,y)$ with $|x|=1$ has the SOP.
Proof.
Theorem 2.2 gives an $L(\mathbb{M})$-formula $\psi(x;y)$ with the SOP, with $|x|=1$. Write $\psi(x,y)$ as $\varphi(x,y,c)$ for some tuple $c$ in $\mathbb{M}$. Then $\varphi(x;y,z)$ has the SOP. Indeed, if
$\psi(\mathbb{M},b_{0})\subsetneq\psi(\mathbb{M},b_{1})\subsetneq\cdots,$
then
$\varphi(\mathbb{M},b_{0},c)\subsetneq\varphi(\mathbb{M},b_{1},c)\subsetneq\cdots\qed$
3 The NIP case
Theorem 3.1.
Suppose $T$ is NIP but has the SOP. Then there is an $L(\mathbb{M})$-formula $\varphi(x,y)$ with the SOP, with $|x|=|y|=1$.
The proof of Theorem 3.1 is really just the standard argument that stability is NSOP plus NIP (see [4, Theorem 12.38] or [7, Theorem 2.67]). Nevertheless, we trace through the details for the sake of completeness.
Proof.
By Theorem 2.2, it suffices to find an SOP $L(\mathbb{M})$-formula $\varphi(x,y)$ with $|y|=1$. Suppose no such formula exists. Then the following holds:
Claim 3.2.
Suppose $C\subseteq\mathbb{M}$ is small, $\{b_{i}\}_{i\in I}$ is a $C$-indiscernible sequence of singletons, $q(x)$ is a partial type over $C$, $\varphi(x,y)$ is an $L(C)$-formula, and $i_{0}<i_{1}$. Then
$\{\varphi(x,b_{i_{0}}),\neg\varphi(x,b_{i_{1}})\}\cup q(x)\text{ is consistent}$
if and only if $\{\neg\varphi(x,b_{i_{0}}),\varphi(x,b_{i_{1}})\}\cup q(x)\text{ is consistent}.$
Proof.
By compactness we can assume $q(x)$ is a single $L(C)$-formula $\psi(x)$. If the claim failed, we would get
either $\varphi(\mathbb{M},b_{i_{0}})\cap\psi(\mathbb{M})\subsetneq\varphi(\mathbb{M},b_{i_{1}})\cap\psi(\mathbb{M})$
or $\varphi(\mathbb{M},b_{i_{0}})\cap\psi(\mathbb{M})\supsetneq\varphi(\mathbb{M},b_{i_{1}})\cap\psi(\mathbb{M}).$
Either way, the formula $\varphi(x,y)\wedge\psi(x)$ has the SOP by indiscernibility. ∎
Since $T$ is unstable, there is an unstable $L$-formula $\psi(x,y)$ with $|y|=1$ [6, Theorem 2.13]. Take an indiscernible sequence $\{(a_{i},b_{i})\}_{i\in\mathbb{Q}}$ such that
$\mathbb{M}\models\psi(a_{i},b_{j})\iff i<j.$
Fix some $n$ and let $[n]=\{1,\ldots,n\}$. For any $S\subseteq[n]$ let $p_{S}(x)$ be the type
$p_{S}(x)=\{\psi(x,b_{i}):i\in[n]\cap S\}\cup\{\neg\psi(x,b_{i}):i\in[n]\setminus S\}$
Claim 3.3.
If $p_{S}$ is consistent and $\sigma$ is a permutation of $[n]$, then $p_{\sigma(S)}$ is consistent.
Proof.
We may assume that $\sigma$ is the permutation $(i~{}i+1)$ swapping $i$ and $i+1$. Let
$q(x)=\{\psi(x,b_{j}):j\in[n]\cap S,~{}j\notin\{i,i+1\}\}\cup\{\neg\psi(x,b_{j}):j\in[n]\setminus S,~{}j\notin\{i,i+1\}\},$
i.e., the part of $p_{S}$ not involving $i$ and $i+1$. We must show
$\{\varphi(x,b_{i}),\neg\varphi(x,b_{i+1})\}\cup q(x)\text{ is consistent iff }\{\neg\varphi(x,b_{i}),\varphi(x,b_{i+1})\}\cup q(x)\text{ is consistent.}$.
This follows by Claim 3.2, applied to the parameter set $C=\{b_{j}:j\in[n]\setminus\{i,i+1\}\}$ and the $C$-indiscernible sequence $\{b_{j}\}_{i-1<j<i+2}$. ∎
Every $S\subseteq[n]$ has the form $\sigma([i])$ for some $i\leq n$ and permutation $\sigma:[n]\to[n]$. Since $p_{[i]}(x)$ is realized by $a_{i+0.5}$, it follows that any $p_{S}$ is consistent. As $n$ and $S$ were arbitrary, this contradicts NIP. ∎
4 Switchboards
4.1 Motivation
We want to produce a theory $T$ such that
  • some formula $\varphi(x,y)$ has the SOP with $|y|=n>1$
  • no $L(M)$-formula $\varphi(x,y)$ with $|y|=1$ has the SOP.
We may as well take $n=2$. Then $\varphi$ determines a partial order $<$ on $M^{2}$ with infinite height. For any fixed $a\in M$, we can restrict this partial order to $\{a\}\times M$. The resulting partial order must have finite height, or some one-variable formula witnesses SOP. Then ($(\{a\}\times M,<)$ must be a finite union of antichains. We may as well require $\{a\}\times M$ to be an antichain. Similarly, we may as well require $M\times\{a\}$ to be an antichain.
So we are now considering the theory of structures $(M,<)$ where $<$ is a partial order on $M^{2}$, such that each set of the form $\{a\}\times M$ or $M\times\{a\}$ is an antichain. The next natural move is to take the model companion, cross our fingers, and hope for everything to work out.
This is the strategy we will follow, but with one further twist: we take the order $<$ on the set of unordered pairs rather than ordered pairs to minimize the number of cases that must be checked in the proof.
4.2 The definitions and goal
If $M$ is a set, let $[M]^{2}$ be the set of 2-element subsets of $M$ (not including singletons). We refer to elements of $[M]^{2}$ as edges, thinking of $M$ as a complete graph.
Definition 4.1.
A switchboard is a structure $(M,<)$, where $<$ is a partial order on $[M]^{2}$, such that $\{x,y\}$ and $\{x,z\}$ are incomparable for any distinct $x,y,z\in M$. In other words, the set of edges incident to $x$ is an antichain, for any $x\in M$. We refer to this condition as the switchboard axiom.
Officially, we regard switchboards as structures in a language $L^{-}$ with a 4-ary relation symbol ${<}(x,y,z,w)$ interpreted as
$x\neq y\wedge z\neq w\wedge\{x,y\}<\{z,w\}.$
Remark 4.2.
A switchboard structure on $M$ is determined by the information of whether $\{x,y\}<\{z,w\}$ holds, for distinct $x,y,z,w\in M$, because $\{x,y\}<\{z,w\}$ can only hold when $x,y,z,w$ are all distinct.
For example, there is a unique switchboard structure on any 3-element set $M$, since we cannot find any four distinct elements $x,y,z,w\in M$.
Theorem 4.3.
  1. 1.
    The theory of switchboards has a model companion $T^{-}$.
  2. 2.
    $T^{-}$ is $\aleph_{0}$-categorical and has the SOP.
  3. 3.
    If $M\models T^{-}$ and $\varphi(x,y)$ is an $L(M)$-formula with $|y|=1$, then $\varphi$ is NSOP.
The proof occupies the rest of this paper.
4.3 Labeled switchboards
Unfortunately, the theory $T^{-}$ fails to have quantifier elimination, because the class of switchboards does not have the amalgamation property (by Remark 4.24 below). We must first work in an expanded language, construct its model companion, then relate it back to the original setting.
Definition 4.4.
A labeled switchboard is a structure $(M,<,\uparrow,\downarrow)$ where
  1. 1.
    $(M,<)$ is a switchboard.
  2. 2.
    $\uparrow$ and $\downarrow$ are binary relations between $M$ and $[M]^{2}$.
  3. 3.
    (Trichotomy axiom) For any $a\in M$ and $\{b,c\}\in[M]^{2}$, exactly one of the following holds:
    • $a\uparrow\{b,c\}$
    • $a\in\{b,c\}$
    • $a\downarrow\{b,c\}$.
  4. 4.
    (Upward axiom) If $a\uparrow\{b,c\}$ and $\{b,c\}<\{b^{\prime},c^{\prime}\}$, then $a\uparrow\{b^{\prime},c^{\prime}\}$.
  5. 5.
    (Downward axiom) If $a\downarrow\{b,c\}$ and $\{b,c\}>\{b^{\prime},c^{\prime}\}$, then $a\downarrow\{b^{\prime},c^{\prime}\}$.
We pronounce $x\uparrow\{y,z\}$ as “$x$ favors $\{y,z\}$” and $x\downarrow\{y,z\}$ as “$x$ disfavors $\{y,z\}$.” Officially, we regard labeled switchboards as structures in a language $L^{+}$ with a 4-ary relation symbol $<$ and two 3-ary relation symbols $\uparrow$ and $\downarrow$. Note that $L^{+}$ expands $L^{-}$.
Remark 4.5.
The last three axioms of labeled switchboards can be understood as saying that for each element $a$, we have a partition of the poset $([M]^{2},<)$ into three sets:
  • An upward-closed set $\{\{x,y\}\in[M]^{2}:a\uparrow\{x,y\}\}$.
  • The antichain $\{\{x,y\}\in[M]^{2}:a\in\{x,y\}\}$.
  • A downward-closed set $\{\{x,y\}\in[M]^{2}:a\downarrow\{x,y\}\}$.
Remark 4.6.
A labeled switchboard structure with underlying set $M$ is determined by the following data:
  • How $\{x,y\}$ and $\{z,w\}$ compare, for distinct $x,y,z,w\in M$.
  • Whether $x\uparrow\{y,z\}$ or $x\downarrow\{y,z\}$ holds, for distinct $x,y,z\in M$.
For example, there are eight different labeled switchboard structures on a three-element set $\{a,b,c\}$, since the relation $<$ between edges can never hold, and then the $\uparrow$ and $\downarrow$ relations can be chosen freely.
Observation 4.7.
Let $P$ be a poset partitioned into three sets $D\sqcup A\sqcup U$, where $A$ is an antichain, $D$ is downward closed, and $U$ is upward closed. If $a\in A$ and $a<x$, then $x\in U$. Indeed,
  • $x\notin A$ since $A$ is an antichain.
  • $x\notin D$, or else $a<x$ implies $a\in D$, contradicting the fact that $a\in A$.
Remark 4.8.
In a labeled switchboard,
$\{a,x\}<\{y,z\}\implies a\uparrow\{y,z\}$
$\{a,x\}>\{y,z\}\implies a\downarrow\{y,z\}.$
For example, the first line follows by applying Observation 4.7 to the partition from Remark 4.5: the element $\{a,x\}$ belongs to the antichain for $a$, so $\{y,z\}$ must be in the upward-closed set for $a$.
Later, in the model companion, it will turn out that
$a\uparrow\{y,z\}\iff\exists x:\{a,x\}<\{y,z\}$
$a\downarrow\{y,z\}\iff\exists x:\{a,x\}>\{y,z\},$
so the two relations $\uparrow$ and $\downarrow$ will be definable from $<$. This will be the bridge between labeled and unlabeled switchboards.
Proposition 4.9.
Every switchboard $(M,<)$ can be expanded to a labeled switchboard $(M,<,\uparrow,\downarrow)$.
Proof.
For distinct $a,x,y\in M$,
  • let $a\uparrow\{x,y\}$ hold if there is $z\in M$ such that $\{a,z\}<\{x,y\}$.
  • let $a\downarrow\{x,y\}$ hold otherwise.
Then $(M,<,\uparrow,\downarrow)$ is a labeled switchboard by Observation 4.10 below. ∎
Observation 4.10.
Let $(P,<)$ be a poset and $A\subseteq P$ be an antichain. If we let
$U$ =$=\{x\in P\mid\exists a\in A:x>a\}$
$D$ $=P\setminus(U\cup A)$
then $U$ is upward-closed, $D$ is downward-closed, and $U\sqcup A\sqcup D$ is a partition of $P$.
Remark 4.11.
The axioms of labeled switchboards are symmetric between $\uparrow$ and $\downarrow$, but the proof of Proposition 4.9 broke this symmetry by treating $\downarrow$ as the default option. This theme will continue in the next section.
4.4 Single-element free amalgamation
In this section, we show that labeled switchboards can be amalgamated, which will help construct the model companion in the next section. Moreover, they can be amalgamated in a specific way (“freely”) which will be useful in the proof that $\varphi(x,y)$ is NSOP whenever $|y|=1$.
First, we reformulate the definition of “labeled switchboard” in a way that is asymmetric and strange, but easier to use for the proof of free amalgamation.
Lemma 4.12.
[…]
  1. (1)
[…]
  2. (2)
    If $\{x,y\}\lhd\{z,w\}$ then $x\lhd\{z,w\}$.
  3. (3)
    If $a\lhd b$, then $b\in[M]^{2}$ (rather than $b\in M$).
  4. (4)
[…]
  5. (5)
[…]
Proof.
First suppose we have a labeled switchboard. Axiom (1) holds either because $<$ is a transitive relation on $[M]^{2}$ or because of the Upward Axiom for labeled switchboards: $x\uparrow\{y,z\}<\{v,w\}\implies x\uparrow\{v,w\}$. Axiom (2) is Remark 4.8. Axiom (3) is obvious. Axiom (4) holds because $<$ is irreflexive. Axiom (5) is part of the Trichotomy Axiom for labeled switchboards.
[…]
$\{x,y\}<\{z,w\}$$\iff\{x,y\}\lhd\{z,w\}$
$x\uparrow\{y,z\}$ $\iff x\lhd\{y,z\}$
[…]
$x\lhd\{y,z\}\lhd\{u,v\}\implies x\lhd\{u,v\},$
[…]
  • $x\uparrow\{u,v\}$. Then $x\lhd\{u,v\}\lhd\{y,z\}$ so $x\lhd\{y,z\}$.
  • $x\in\{u,v\}$. Then $x\in\{u,v\}\lhd\{y,z\}$ so $x\lhd\{y,z\}$ by Axiom (2).
Either way, $x\lhd\{y,z\}$, so $x\uparrow\{y,z\}$, contradicting the fact that $x\downarrow\{y,z\}$. Thus all the axioms of labeled switchboards hold. ∎
Definition 4.13.
[…]
Lemma 4.14.
Let $X_{1},X_{2}$ be two sets. Let $<_{i}$ be a transitive relation on $X_{i}$ for $i=1,2$, such that $<_{1}$ and $<_{2}$ have the same restriction to $X_{1}\cap X_{2}$. Let $<$ be the transitive closure of the union of $<_{1}$ and $<_{2}$.
  1. 1.
    $<$ extends $<_{i}$ on $X_{i}$.
  2. 2.
    If $a\in X_{1}\setminus X_{2}$ and $b\in X_{2}\setminus X_{1}$, then $a<b$ holds if and only if there is $c\in X_{1}\cap X_{2}$ such that $a<_{1}c<_{2}b$.
  3. 3.
    If $a\in X_{2}\setminus X_{1}$ and $b\in X_{1}\setminus X_{2}$, then $a<b$ holds if and only if there is $c\in X_{1}\cap X_{2}$ such that $a<_{2}c<_{1}b$.
This is well-known, but we include the proof for completeness.
Proof.
It suffices to prove the following:
Claim 4.15.
If $a<b$, then one of the following holds:
  1. 1.
    $a<_{i}b$ for $i=1$ or $2$. In particular, $\{a,b\}\subseteq X_{i}$.
  2. 2.
    $a\in X_{1}\setminus X_{2}$ and $b\in X_{2}\setminus X_{1}$ and $a<_{1}c<_{2}b$ for some $c\in X_{1}\cap X_{2}$.
  3. 3.
    $a\in X_{2}\setminus X_{1}$ and $b\in X_{1}\setminus X_{2}$ and $a<_{2}c<_{1}b$ for some $c\in X_{1}\cap X_{2}$.
Let $<_{0}$ be the union of $<_{1}$ and $<_{2}$. By construction of the transitive closure, there is a sequence
$a=z_{0}<_{0}z_{1}<_{0}\cdots<_{0}z_{n}=b$
with $n>0$. Take such a sequence with $n$ minimal. If $z_{j}\in X_{1}\setminus X_{2}$ for some $0<j<n$, then we must have $z_{j-1}<_{1}z_{j}<_{1}z_{j+1}$ (since $z_{j}$ cannot satisfy $<_{2}$). By transitivity of $<_{1}$, we can drop $z_{j}$ from the list, contradicting minimality. Similarly, $z_{j}$ cannot be in $X_{2}\setminus X_{1}$ for $0<j<n$. Therefore, $z_{j}\in X_{1}\cap X_{2}$ for each $0<j<n$. Break into cases depending on $n$.
  • $n=1$. Then $a<_{0}b$, so we are in Case 1 of the Claim.
  • $n=2$. Then $a<_{0}z<_{0}b$ for some $z\in X_{1}\cap X_{2}$. If $a,b$ are both in $X_{1}$, then $a<_{1}z<_{1}b$, contradicting minimality. So one of $a$ and $b$ is in $X_{2}\setminus X_{1}$. Similarly, one is in $X_{1}\setminus X_{2}$. Then we are Case 2 or 3 of the Claim.
  • $n>2$. Take $i$ such that $a\in X_{i}$. Then $a,z_{1},z_{2}$ are in $X_{i}$, so $a<_{i}z_{1}<_{i}z_{2}$. Then $z_{1}$ could be dropped from the list, contradicting minimality. ∎
Definition 4.16.
Let $M$ be a labeled switchboard. Let $S$ be a subset and let $a_{1},a_{2}$ be two distinct elements of $M$. Then $a_{1}$ and $a_{2}$ are freely amalgamated over $S$ if the following holds:
  1. (i)
    If $x,y\in S$, then $\{a_{1},x\}<\{a_{2},y\}$ if and only if there is $\{p,q\}\in[S]^{2}$ with $\{a_{1},x\}<\{p,q\}$ and $\{p,q\}<\{a_{2},y\}$.
  2. (ii)
    If $x,y\in S$, then $\{a_{2},y\}<\{a_{1},x\}$ if and only if there is $\{p,q\}\in[S]^{2}$ with $\{a_{2},y\}<\{p,q\}$ and $\{p,q\}<\{a_{1},x\}$.
  3. (iii)
    The new edge $\{a_{1},a_{2}\}$ is incomparable to every other element of $[S\cup\{a_{1},a_{2}\}]^{2}$.
  4. (iv)
    If $x\in S$, then $x\downarrow\{a_{1},a_{2}\}$.
  5. (v)
    If $x\in S$, then $a_{1}\uparrow\{a_{2},x\}$ if and only if there is $\{p,q\}\in[S]^{2}$ such that $a_{1}\uparrow\{p,q\}$ and $\{p,q\}<\{a_{2},x\}$.
  6. (vi)
    If $x\in S$, then $a_{2}\uparrow\{a_{1},x\}$ if and only if there is $\{p,q\}\in[S]^{2}$ such that $a_{2}\uparrow\{p,q\}$ and $\{p,q\}<\{a_{1},x\}$.
Lemma 4.17.
Suppose $S,A_{1},A_{2}$ are labeled switchboards such that
  • $A_{1}$ and $A_{2}$ extend $S$.
  • $A_{i}=S\cup\{a_{i}\}$ for $i=1,2$, for two distinct elements $a_{1},a_{2}\notin S$.
Let $M=A_{1}\cup A_{2}=S\cup\{a_{1},a_{2}\}$. Then we can make $M$ into a labeled switchboard extending $A_{1}$ and $A_{2}$, in which $a_{1}$ and $a_{2}$ are freely amalgamated over $S$.
Proof.
[…]
Claim 4.18.
[…]
Note that $(A_{1}\cup[A_{1}]^{2})\cap(A_{2}\cup[A_{2}]^{2})=S\cup[S]^{2}$. By Lemma 4.14, the following claims hold:
Claim 4.19.
[…]
Claim 4.20.
If $x$ is in $(A_{i}\cup[A_{i}]^{2})\setminus(S\cup[S]^{2})$, and $y$ is in $(A_{j}\cup[A_{j}]^{2})\setminus(S\cup[S]^{2})$ for $i\neq j$, then $x\lhd y$ holds if and only if there is $z\in S\cup[S]^{2}$ such that $x\lhd_{i}z\lhd_{j}y$.
[…]
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
    If $a\lhd b$, then $b\in[M]^{2}$: By construction of the transitive closure there is $b^{\prime}$ and $i$ such that $b^{\prime}\lhd_{i}b$. Since $\lhd_{i}$ itself satisfies Axiom (3) of the triangle axioms, $b$ is in $[M]^{2}$ rather than $M$.
  4. 4.
[…]
  5. 5.
[…]
    • If $\{x,y\}\subseteq A_{i}$ for some $i$, then $x\lhd_{i}\{x,y\}$ by Claim 4.19, contradicting the fact that $\lhd_{i}$ itself satisfies Axiom (5) of the triangle axioms.
    • Otherwise, $\{x,y\}=\{a_{1},a_{2}\}$, and then $x\lhd\{a_{1},a_{2}\}$ contradicts Claim 4.18.
Finally, we check that Conditions (i)(vi) in Definition 4.16 hold.
  1. (i)
    This says that $\{a_{1},x\}\lhd\{a_{2},y\}$ if and only if there is $p\in[S]^{2}$ such that $\{a_{1},x\}\lhd_{1}p\lhd_{2}\{a_{2},y\}$. This is an instance of Claim 4.20, since $\{a_{1},x\}$ is from $[A_{1}]^{2}\setminus[S]^{2}$ and $\{a_{2},y\}$ is from $[A_{2}]^{2}\setminus[S]^{2}$.
  2. (ii)
    Similar.
  3. (iii)
[…]
  4. (iv)
[…]
  5. (v)
    This says that $a_{1}\lhd\{a_{2},x\}$ if and only if there is $p\in[S]^{2}$ such that $a_{1}\lhd_{1}p\lhd_{2}\{a_{2},x\}$. This is another instance of Claim 4.20.
  6. (vi)
    Similar. ∎
The reader can now safely forget about triangle relations.
4.5 The model companion $T$
Proposition 4.21.
The class of finite labeled switchboards is a Fraïsse class.
Proof.
Hereditary property:
Clear, since the axioms are $\forall$-sentences.
Amalgamation property:
Suppose we are amalgamating this picture:
${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B}$${C.}$
Since we are in a purely relational language, we can use induction on the size of $|B\setminus A|$ and $|C\setminus A|$ to reduce to the case where $|B\setminus A|=|C\setminus A|=1$. This case is handled by Lemma 4.17.
Joint embedding property:
Use the amalgamation property over the empty labeled switchboard. ∎
Let $M_{0}$ be the Fraïsse limit of finite labeled switchboards. Let $T^{+}$ be the complete theory of $M_{0}$. By general machinery, the following holds:
  • $T^{+}$ has quantifier elimination.
  • $T^{+}$ is complete and countably categorical.
  • $T^{+}$ is the model companion of labeled switchboards.
  • The class of labeled switchboards has the amalgamation property.
4.6 Comparison to unlabeled switchboards
Lemma 4.22.
If $x,y,z$ are distinct elements of a model $M$ of $T^{+}$, then
$x\uparrow\{y,z\}\iff\exists w\neq x:\{x,w\}<\{y,z\}$
$x\downarrow\{y,z\}\iff\exists w\neq x:\{x,w\}>\{y,z\}.$
Proof.
The right-to-left direction was a consequence of the axioms of labeled switchboards; see Remark 4.8. For the left-to-right direction, suppose $x\uparrow\{y,z\}$. We need to find $w\in M$ such that $\{x,w\}<\{y,z\}$. Since the model $M$ is existentially closed among labeled switchboards, it suffices to instead find $w$ in an extension $N\supseteq M$. Let $w$ be a point outside $M$. Make the four-element set $\{x,y,z,w\}$ into a switchboard by making $\{x,w\}<\{y,z\}$ and no other relations hold, so that the poset [$[\{x,y,z,w\}]^{2}$ looks like this:
${\{y,z\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\{x,w\}}$${\{x,y\}}$${\{x,z\}}$${\{y,w\}}$${\{z,w\}}$
Expand $\{x,y,z,w\}$ to a labeled switchboard by making
  • $x\uparrow\{y,z\}$ and $w\uparrow\{y,z\}$ and $y\downarrow\{x,w\}$ and $z\downarrow\{x,w\}$.
  • $y\uparrow\{x,z\}$ if and only if it holds in $M$.
  • $z\uparrow\{x,y\}$ if and only if it holds in $M$.
  • All other information chosen randomly.
The reader can verify that the upward and downward axioms (i.e., Axioms (4) and (5) in Definition 4.4) hold [^1] Note that these axioms only need to be checked relative to the inequality $\{x,w\}<\{y,z\}$, as it is the sole inequality that holds..
Use the amalgmation property for labeled switchboards to amalgamate $\{x,y,z,w\}$ and $M$ together over $\{x,y,z\}$:
${\{x,y,z\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\{x,y,z,w\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N}$
Then we get a bigger labeled switchboard $N$ extending $M$, and containing an element $w$ such that $\{x,w\}<\{y,z\}$. ∎
So in the theory $T^{+}$, the two relations $\uparrow$ and $\downarrow$ are definable from the relation $<$. Let $T^{-}$ be the reduct of $T^{+}$ to the language without $\uparrow$ and $\downarrow$. Then $T^{+}$ is a definitional expansion of $T^{-}$. So $T^{-}$ has the same semantic properties as $T^{+}$. For example, $T^{-}$ is countably categorical and complete.
Theorem 4.23.
$T^{-}$ is the model companion of (unlabeled) switchboards.
Proof.
It suffices to check the following three claims:
  1. 1.
    Every switchboard embeds into a model of $T^{-}$: first use Proposition 4.9 to expand $M$ to a labeled switchboard, then embed it into a model of $T^{+}$, then take the reduct to the language $L^{-}$.
  2. 2.
    Every model of $T^{-}$ is a switchboard: clear.
  3. 3.
    $T^{-}$ is model complete: this holds because $T^{-}$ has quantifier elimination after adding the two symbols $\uparrow$,$\downarrow$ , ↓, and these symbols are both existentially definable and universally definable. The existentialy definability is Lemma 4.22. The universal definability holds because $\uparrow$ and $\downarrow$ are essentially each other’s complements, by the trichotomy axiom. ∎
Remark 4.24.
The theory $T^{-}$ does not have quantifier elimination. Otherwise, every switchboard $M$ would have a unique extension to a labeled switchboard, contradicting the fact that the unlabeled switchboard with three elements has eight extensions to a labeled switchboard (Remarks 4.2 and 4.6). It follows that the class of switchboards does not have the amalgamation property. Concrete failures of the AP can be extracted from the proof of Lemma 4.22.
4.7 The theory has the SOP
Make $\mathbb{Z}$ into a switchboard by ordering
<$\cdots<\{0,1\}<\{2,3\}<\{4,5\}<\cdots$
and making everything else in $[\mathbb{Z}]^{2}$ be incomparable. Embed $\mathbb{Z}$ into a model $(M,<)$ of $T^{-}$. Then the poset $([M]^{2},<)$ has an infinite chain, inherited from $([\mathbb{Z}]^{2},<)$. Therefore $T^{-}$ and the equivalent theory $T^{+}$ have the SOP.
4.8 One-variable formulas
Work in a monster model $\mathbb{M}$ of $T^{+}$.
Definition 4.25.
Let $B\subseteq\mathbb{M}$ be small, and let $a_{1},a_{2}\in\mathbb{M}^{1}$ be two elements with $a_{1}\equiv_{B}a_{2}$.
  1. 1.
    $\operatorname{tp}(a_{1},a_{2}/\mathbb{M})$ is half-symmetric if
    $\{a_{1},b\}<\{a_{2},c\}\iff\{a_{2},b\}<\{a_{1},c\}$
    for any $\{b,c\}\in[B]^{2}$.
  2. 2.
    $\operatorname{tp}(a_{1},a_{2}/\mathbb{M})$ is symmetric if it is half-symmetric, and
    $a_{1}\uparrow\{a_{2},b\}\iff a_{2}\uparrow\{a_{1},b\}$
    $a_{1}\downarrow\{a_{2},b\}\iff a_{2}\downarrow\{a_{1},b\}$
    for any $b\in B$.
Remark 4.26.
If $a_{1},a_{2}$ are singletons and $a_{1}\equiv_{B}a_{2}$, then $\operatorname{tp}(a_{1},a_{2}/B)$ is symmetric if and only if $a_{1}a_{2}\equiv_{B}a_{2}a_{1}$. This can be seen by quantifier elimination—the only atomic formulas in $\operatorname{tp}(a_{1},a_{2}/B)$ beyond those in $\operatorname{tp}(a_{1}/B)\cup\operatorname{tp}(a_{2}/B)$ are the formulas
$\{x_{1},b\}<\{x_{2},c\},\text{ etc.}$.
$x_{1}\uparrow\{x_{2},b\},\text{ etc.}$.
appearing in Definition 4.25, and intrinsically symmetric formulas like
$\{x_{1},x_{2}\}<\{b,c\},\text{ etc.}$.
$b\uparrow\{x_{1},x_{2}\},\text{ etc.}$.
Definition 4.27.
Let $a_{1},a_{2}$ be distinct singletons in $\mathbb{M}\setminus B$. Then $\operatorname{tp}(a_{1},a_{2}/B)$ is distinguished if the following two conditions both hold:
  • For any $b,c\in B$, if $\{a_{1},b\}>\{a_{2},c\}$, then there is $\{u,v\}\in[B]^{2}$ such that $\{a_{1},b\}>\{u,v\}>\{a_{2},c\}$.
  • For any $b,c\in B$, if $\{a_{1},b\}<\{a_{2},c\}$, then there is $\{u,v\}\in[B]^{2}$ such that $\{a_{1},b\}<\{u,v\}<\{a_{2},c\}$.
Lemma 4.28.
If $a_{1}\equiv_{B}a_{2}$ and $\operatorname{tp}(a_{1},a_{2}/B)$ is distinguished, then $\operatorname{tp}(a_{1},a_{2}/B)$ is half-symmetric.
Proof.
We must show that for any $b,c\in B$,
$\{a_{1},b\}<\{a_{2},c\}\iff\{a_{2},b\}<\{a_{1},c\}.$
We prove the $\Rightarrow$ direction; the $\Leftarrow$ direction is similar. Suppose $\{a_{1},b\}<\{a_{2},c\}$. By definition of “distinguished”, there is $\{u,v\}\in[B]^{2}$ such that $\{a_{1},b\}<\{u,v\}<\{a_{2},c\}$. Since $a_{1}\equiv_{B}a_{2}$, and $\{b,c,u,v\}\subseteq B$, we have
$\{a_{1},b\}<\{u,v\}\implies\{a_{2},b\}<\{u,v\}$
$\{u,v\}<\{a_{2},c\}\implies\{u,v\}<\{a_{1},c\}.$
Then
$\{a_{2},b\}<\{u,v\}<\{a_{1},c\},$
so $\{a_{2},b\}<\{a_{1},c\}$ as desired. ∎
Lemma 4.29.
Let $B$ be a small set and let $c_{1},c_{2},c_{3}$ be distinct elements of $\mathbb{M}\setminus B$, all realizing the same type over $B$. Suppose
  • $c_{1}$ and $c_{3}$ are freely amalgated over $Bc_{2}$ in the sense of Definition 4.16.
  • $\operatorname{tp}(c_{1},c_{2}/B)$ and $\operatorname{tp}(c_{1},c_{3}/B)$ and $\operatorname{tp}(c_{2},c_{3}/B)$ are distinguished, hence half-symmetric.
Then $\operatorname{tp}(c_{1},c_{3}/B)$ is symmetric.
Proof.
By assumption, $\operatorname{tp}(c_{1},c_{3}/B)$ is half-symmetric, so it remains to prove that
$c_{1}\uparrow\{c_{3},b\}\iff c_{3}\uparrow\{c_{1},b\},$
for $b\in B$. By symmetry, it suffices to prove the $\Rightarrow$ direction. Suppose $c_{1}\uparrow\{c_{3},b\}$. By the definition of free amalgamation (specifically, Condition (v) in Definition 4.16), there is $\{u_{0},v_{0}\}\subseteq Bc_{2}$ such that
$c_{1}\uparrow\{u_{0},v_{0}\}\text{ and }\{u_{0},v_{0}\}<\{c_{3},b\}.$
Claim 4.30.
There is $\{u,v\}\subseteq B$ such that
$c_{1}\uparrow\{u,v\}\text{ and }\{u,v\}<\{c_{3},b\}.$
Proof.
If $\{u_{0},v_{0}\}\subseteq B$, take $\{u,v\}=\{u_{0},v_{0}\}$. Otherwise, $c_{2}\in\{u_{0},v_{0}\}$, so $\{u_{0},v_{0}\}=\{c_{2},b^{\prime}\}$ for some $b^{\prime}\in B$. Then
$c_{1}\uparrow\{c_{2},b^{\prime}\}\text{ and }\{c_{2},b^{\prime}\}<\{c_{3},b\}.$
Since $\operatorname{tp}(c_{2},c_{3}/B)$ is distinguished, there is some $\{u,v\}\in[B]^{2}$ such that
$\{c_{2},b^{\prime}\}<\{u,v\}<\{c_{3},b\}.$
By the Upward Axiom of labeled switchboards,
$c_{1}\uparrow\{c_{2},b^{\prime}\}<\{u,v\}\implies c_{1}\uparrow\{u,v\}.\qed$
Since $c_{1}\equiv_{B}c_{3}$ and $\{u,v,b\}\subseteq B$,
$c_{1}\uparrow\{u,v\}\implies c_{3}\uparrow\{u,v\}$
$\{u,v\}<\{c_{3},b\}\implies\{u,v\}<\{c_{1},b\}.$
Finally,
$c_{3}\uparrow\{u,v\}<\{c_{1},b\}\implies c_{3}\uparrow\{c_{1},b\}$
by the Upward Axiom. ∎
Lemma 4.31.
Let $B$ be a small set and let $a_{1},a_{2}$ be two elements of $\mathbb{M}\setminus B$. Then there is $\sigma\in\operatorname{Aut}(\mathbb{M}/B)$ such that $a_{1}$ and $\sigma(a_{2})$ are freely amalgamated over $B$.
Proof.
This follows formally from Lemma 4.17 and the fact that $T^{+}$ is the model completion of labeled switchboards. More precisely, take two distinct elements $c_{1},c_{2}$ outside $B$, and make $B\cup\{c_{i}\}$ into a labeled switchboard isomorphic to $B\cup\{a_{i}\}$. Use Lemma 4.17 to make $B\cup\{c_{1},c_{2}\}$ into a labeled switchboard in which $c_{1}$ and $c_{2}$ are freely amalgamated. Use quantifier elimination and the fact that $\mathbb{M}$ is a monster model to embed $B\cup\{c_{1},c_{2}\}$ into $\mathbb{M}$ over $B$. The images of $c_{1}$ and $c_{2}$ give two elements $e_{1},e_{2}\in\mathbb{M}$ such that $e_{i}\equiv_{B}c_{i}\equiv_{B}a_{i}$ for $i=1,2$, and $e_{1}$ and $e_{2}$ are freely amalgated over $B$. Use a further automorphism to move $e_{1}$ to $a_{1}$. ∎
Proposition 4.32.
Let $B$ be a finite subset of $\mathbb{M}$, let $p(x)$ be a complete 1-type over $B$ not realized in $B$, and let $q(x,y)$ be a complete 2-type over $B$ extending $p(x)\cup p(y)$.
  1. 1.
    There exists a sequence $c_{0},c_{1},c_{2},\ldots$ of realizations of $p$, such that
    • $c_{i}c_{i+1}$ realizes $q$ for each $i$.
    • For $i\geq 2$, $c_{i}$ and $c_{0}$ are freely amalgamated over $Bc_{i-1}$.
For the remaining two points fix a sequence $c_{0},c_{1},c_{2},\ldots$ as in the the previous point, and assume that the elements $c_{0},c_{1},c_{2},\ldots$ are pairwise distinct.
  1. 2.
    If $i\geq|B|$, then $\operatorname{tp}(c_{0},c_{i}/B)$ is distinguished, hence half-symmetric.
  2. 3.
    If $i>|B|$, and $q$ is distinguished, then $\operatorname{tp}(c_{0},c_{i}/B)$ is symmetric.
Proof.
  1. 1.
    Take $c_{0}c_{1}$ to be any realization of $p$. For $i\geq 2$, take $c_{i}$ such that $c_{i-1}c_{i}\models q$, then use Lemma 4.31 to move $c_{i}$ by an automorphism over $Bc_{i-1}$ to make $c_{i}$ and $c_{0}$ be freely amalgated over $Bc_{i-1}$.
  2. 2.
    First we prove the following claim:
    Claim 4.33.
    Suppose $i>0$ and $b_{i},b_{0}$ are elements of $B$ such that $\{c_{i},b_{i}\}>\{c_{0},b_{0}\}$. Then one of the following happens:
    1. (a)
      There is $\{u,v\}\in[B]^{2}$ such that $\{c_{i},b_{i}\}>\{u,v\}>\{c_{0},b_{0}\}$.
    2. (b)
      There are $b_{1},\ldots,b_{i-1}\in B$ such that
      $\{c_{i},b_{i}\}>\{c_{i-1},b_{i-1}\}>\cdots>\{c_{0},b_{0}\}.$
    Proof.
    Proceed by induction on $i$. When $i=1$, Case (b) holds vacuously. Suppose $i>1$. Then $c_{i}$ and $c_{0}$ are freely amalgamated over $Bc_{i-1}$. Because of the definition of free amalgamation (i.e., Conditions (i), (ii) of Definition 4.16), there must be $\{u,v\}\in[Bc_{i-1}]^{2}$ such that $\{c_{i},b_{i}\}>\{u,v\}>\{c_{0},b_{0}\}$. If $\{u,v\}\subseteq B$ then we are in Case (a). Otherwise, $c_{i-1}\in\{u,v\}$, so $\{u,v\}=\{c_{i-1},b_{i-1}\}$ for some $b_{i-1}\in B$. Then $\{c_{i-1},b_{i-1}\}>\{c_{0},b_{0}\}$, so by induction, one of the following holds:
    • There are $\{u,v\}\subseteq B$ such that $\{c_{i-1},b_{i-1}\}>\{u,v\}>\{c_{0},b_{0}\}$. Then
      $\{c_{i},b_{i}\}>\{c_{i-1},b_{i-1}\}>\{u,v\}>\{c_{0},b_{0}\}$
      and we are in Case (a).
    • There are $b_{i-2},\ldots,b_{1}$ such that
      $\{c_{i},b_{i}\}>\{c_{i-1},b_{i-1}\}>\{c_{i-2},b_{i-2}\}>\cdots>\{c_{0},b_{0}\}.$
      Then we are in Case (b). ∎
    Now suppose $i\geq|B|$. We must show that $\operatorname{tp}(c_{0},c_{i}/B)$ is distinguished. There are two points to check in Definition 4.27. We check the second point; the first is similar222…using a variant of the claim where we replace $>$ with $<$. The proof is identical, but we can’t just say it works “by symmetry” since the definition of free amalgamation broke the symmetry between up and down in the poset. Suppose $b_{0},b_{i}\in B$ are such that
    $\{c_{i},b_{i}\}>\{c_{0},b_{0}\}.$
    We must find $\{u,v\}\in[B]^{2}$ such that
    $\{c_{i},b_{i}\}>\{u,v\}>\{c_{0},b_{0}\}.$
    Otherwise, the Claim gives $b_{1},b_{2},\ldots,b_{i-1}\in B$ such that
    $\{c_{i},b_{i}\}>\{c_{i-1},b_{i-1}\}>\cdots>\{c_{0},b_{0}\}.$
    By the pigeonhole principle, there are $j_{1}<j_{2}\leq i$ with $b_{j_{1}}=b_{j_{2}}=:b$. Then $\{c_{j_{1}},b\}>\{c_{j_{2}},b\}$, which contradicts the axioms of switchboards.
  3. 3.
    By the previous point and the assumption, $\operatorname{tp}(c_{0},c_{i-1}/B)$ and $\operatorname{tp}(c_{i-1}c_{i}/B)=q$ are both distinguished. By the previous point, $\operatorname{tp}(c_{0},c_{i}/B)$ is distinguished. By Lemma 4.29, applied to the three elements $c_{0},c_{i-1},c_{i}$, we see that $\operatorname{tp}(c_{0},c_{i}/B)$ is symmetric. ∎
Theorem 4.34.
Suppose $M$ is a model of $T^{+}$ and $\varphi(x,y)$ is an $L(M)$-formula with $|y|=1$. Then $\varphi(x,y)$ is NSOP.
Proof.
We may assume $M$ is a monster model $\mathbb{M}$. Take a finite set $B$ containing all the parameters in $\varphi(x,y)$, so that $\varphi$ is an $L(B)$-formula. Take a $B$-indiscernible sequence $a_{0},a_{1},a_{2},\ldots$ in $\mathbb{M}^{1}$ such that
$\varphi(\mathbb{M},a_{0})\subsetneq\varphi(\mathbb{M},a_{1})\subsetneq\cdots$
Let $p=\operatorname{tp}(a_{i}/B)$ for any $i$. Let $q=\operatorname{tp}(a_{i},a_{j}/B)$ for any $i<j$. If $(b,c)\models q$, then $b,c\models p$, and $\varphi(\mathbb{M},b)\subsetneq\varphi(\mathbb{M},c)$.
Let $c_{0},c_{1},c_{2},\ldots$ be a sequence as in Proposition 4.32, with respect to our chosen $p$ and $q$. In particular, $c_{i}c_{i+1}$ realizes $q$ for each $i$, so
$\varphi(\mathbb{M},c_{0})\subsetneq\varphi(\mathbb{M},c_{1})\subsetneq\cdots$
which implies that the $c_{i}$ are pairwise distinct. Take $n=|B|$. By part (2) of Proposition 4.32, $q^{\prime}=\operatorname{tp}(c_{0},c_{n}/B)$ is distinguished. Note that if $(b,c)\models q^{\prime}$, then $\varphi(\mathbb{M},b)\subsetneq\varphi(\mathbb{M},c)$, because $\varphi(\mathbb{M},c_{0})\subsetneq\varphi(\mathbb{M},c_{n})$.
Let $c^{\prime}_{0},c^{\prime}_{1},c^{\prime}_{2},\ldots$ be a sequence as in Proposition 4.32, with respect to $p$ and $q^{\prime}$. In particular, $c^{\prime}_{i}c^{\prime}_{i+1}$ realizes $q^{\prime}$ for each $i$, so
$\varphi(\mathbb{M},c^{\prime}_{0})\subsetneq\varphi(\mathbb{M},c^{\prime}_{1})\subsetneq\cdots$
which implies that the $c^{\prime}_{i}$ are pairwise distinct. Take $n=|B|+1$. By part (3) of Proposition 4.32, $\operatorname{tp}(c^{\prime}_{0},c^{\prime}_{n}/B)$ is symmetric. Then
$\varphi(\mathbb{M},c^{\prime}_{0})\subsetneq\varphi(\mathbb{M},c^{\prime}_{n})\implies\varphi(\mathbb{M},c^{\prime}_{n})\subsetneq\varphi(\mathbb{M},c^{\prime}_{0}),$
a contradiction. ∎
Acknowledgments.
The author was supported by the Ministry of Education of China (Grant No. 22JJD110002). Nick Ramsey encouraged the author to write up this note.